# Overcoming Common Flaws in the Evaluation of Selective Classification Systems

### v1

@@ 1 Introduction
Several current approaches attempt to address this need for multi-threshold metrics in SC. The Area Under the Risk Coverage curve (AURC) is the most prevalent of these Geifman et al. [2018]. However, we demonstrate in this work that AURC has significant limitations as it fails to adequately translate the risk from specific working points into a meaningful aggregated evaluation score. This shortcoming hampers the ability to benchmark and improve SC methodologies effectively.
@@
Proposal of AUGRC: We introduce the Area Under the Generalized Risk Coverage (AUGRC) curve, a new metric designed to overcome the flaws of current multi-threshold metrics for SC. AUGRC meets all five requirements, providing a comprehensive and interpretable measure of SC system performance.
Empirical Validation: We empirically demonstrate the relevance and effectiveness of AUGRC through a comprehensive benchmark spanning 6 datasets and 13 confidence scoring functions.
@@
Figure 1: The AUGRC metric based on Generalized Risk overcomes common flaws in current evaluation of Selective classification (SC). a) Refined task definition for SC. Analogously to standard classification, we distinguish between holistic evaluation for method development and benchmarking using multi-threshold metrics versus evaluation of specific application scenarios at pre-determined working points. The current most prevalent multi-threshold metric in SC, $\mathrm{AURC}$, is based on Selective Risk, a concept for working point evaluation that is not suitable for aggregation over rejection thresholds (red arrow). To fill this gap, we formulate the new concept of Generalized Risk and a corresponding metric, $\mathrm{AUGRC}$ (green arrow). b) We formalize our perspective on SC evaluation by identifying five key requirements for multi-threshold metrics and analyze how previous metrics fail to fulfill them. Abbreviations, CSF: Confidence Scoring Function.
@@ 2 Refined Task Formulation
SC systems consist of a classifier $m:\mathcal{X}\rightarrow\mathcal{Y}$, which outputs a prediction and a CSF $g:\mathcal{X}\rightarrow\mathbb{R}$, which outputs a confidence score associated with the prediction. Assuming a supervised training setup, let $\{(x_{i},y_{i})\}_{i=1}^{N}$ be a dataset containing $N$ independent samples from the source distribution $P(X,Y)$ over $\mathcal{X}\times\mathcal{Y}$. Given a rejection threshold $\tau$, the model prediction is accepted only if the corresponding score is larger than $\tau$:
@@ 2.1 Evaluating SC systems in applied settings
Current SC evaluation commonly reports the selective risk of the system at a given coverage (“risk@coverage”), which implies a pre-determined cutoff on the risk or coverage (e.g. "maximum risk of 5%") leading to a working point of the system defined by a rejection threshold $\tau$. The selective risk is defined as:
@@
This formulation only considers accepted predictions with $c>\tau$. For a binary failure error, this risk is an empirical estimator for the conditional probability $P(Y_{\text{f}}=1|g(x)\geq\tau)$. Thus, this metric effectively communicates the risk of a "silent" failure for a specific prediction of classifier $m$ that has been accepted ($c>\tau$) at a pre-defined threshold $\tau$. This information may be useful in applied medical settings, e.g. to communicate the risk of misclassification for a patient p whose associated prediction has been accepted by a given SC system ($c_{p}>\tau$).
@@ 2.2 Evaluating SC systems for method development and benchmarking
Evaluating the general performance of a SC system requires a multi-threshold metric for a more comprehensive view compared to a working point analysis based on a fixed singular rejection threshold, analogous to binary classification tasks, where the Area Under the Receiver Operating Characteristic (AUROC) is used to assess the entire system performance across multiple classifier thresholds. However, the working point analysis can not be simply translated to multi-threshold aggregation, as it is based on Selective Risk, which only considers the risk w.r.t accepted predictions, assuming that a specific selection has already occurred ($P(Y_{\text{f}}=1\mid g(x)\geq\tau)$). This evaluation ignores rejected cases and thus contradicts the holistic assessment, where we are interested in the general risk for any prediction causing a silent failure independent of the future selection decision of individual predictions. This holistic assessment is reflected by the joint probability $P(Y_{\text{f}}=1,g(x)\geq\tau)$, which reflects the risk of silent failure for any prediction processed by the system including cases that are potentially rejected in the future.
@@ 2.3 Requirements for Selective Classification multi-threshold metrics
Based on the refined task definition given above, we can formulate concrete five requirements R1-R5 for multi-threshold metrics in SC:
R1: Task Completeness. The general goal in SC is to prevent silent failures either by preventing failures in the first place (via classifier performance), or by detecting the remaining failures (via CSF ranking quality). As argued in Jäger et al. [2023], it is crucial to jointly evaluate both aspects of an SC system, since the choice of CSF generally affects the underlying classifier. R2: Monotonicity. The metric should be monotonic w.r.t both evaluated factors stated in R1, i.e. improving on one of the two factors while keeping the other one fixed results in an improved metric value. Note that R2 does not make assumptions about how the two factors are combined, but it represents a minimum requirement for meaningful comparison and optimization of the metric. R3: Ranking Interpretability. Interpretability is a crucial component of a metric [Maier-Hein et al., 2022]. AUROC is the de facto standard ranking metric for binary classification tasks, providing an intuitive assessment of ranking quality which is proportional to the number of permutations needed for establishing an optimal ranking. This can also be interpreted as the "probability of a positive sample having a higher schore than a negative one". Ideally, the SC metric should follow this intuitive assessment of rankings. R4: CSF Flexibility. The metric should be applicable to arbitrary choices of CSFs. This includes external CSFs, i.e. scores that are not based directly on the classifier output. R5: Error Flexibility. Current SC literature largely focuses on 0/1 error (i.e. $1-\text{accuracy}$) in their risk computation. However, in many real-world scenarios, accuracy is not an adequate classification metric, such as in the presence of class imbalance and for pixel-level classification. Thus, the SC metric should be flexible w.r.t the choice of error function.
@@ 2.4 Current multi-threshold metrics in SC do not fulfill requirements R1-R5
This integral effectively aggregates the Selective Risk (Equation 2) over all fractions of accepted predictions (i.e. coverage). However, as discussed in Section 2.2, the Selective Risk is not suitable for aggregation over rejection thresholds to holistically assess a SC system. This inadequacy effectively leads to an excessive over-weighting of high-confidence failures in $\mathrm{AURC}$ and in return to breaching the requirements of monotonicity (R2) and ranking interpretability (R3). We provide empirical evidence for these shortcomings on toy data (Section 3) and real-world data (Section 4.2). Further, current SC literature usually employs the 1/0 error function, even though the related Accuracy metric is well known to be an unsuitable classification performance measure for many applications. For example, Balanced Accuracy is used to address class imbalance and metrics such as the Dice Score are employed for segmentation. Moving beyond the 1/0 error yields a higher variance in distinct error values and may thus amplify the shortcoming of over-weighting high-confidence failures. As more sophisticated error functions being are an important direction of future SC research, it is crucial to ensure the compatibility of metrics in SC evaluation.
e-AURC: Geifman et al. [2018] further introduce the $\mathrm{e\text{-}AURC}$ as the difference to the $\mathrm{AURC}$ of an optimal CSF (denoted as $\mathrm{AURC}$ )
@@ 3 Area under the Generalized Risk Coverage Curve
This metric evaluates SC models in terms of the expected risk of silent failures across working points and thus provides a practical measurement that is directly interpretable. It is bounded to the $[0,\nicefrac{{1}}{{2}}]$ interval, whereby lower $\mathrm{AUGRC}$ corresponds to better SC performance. The $\mathrm{AUGRC}$ is not subject to the shortcomings of the $\mathrm{AURC}$, and we can derive a direct relationship to $\mathrm{AUROC_{f}}$ and $\mathrm{acc}$ (the derivation and visualizations are shown in Equation 8 and Figure 5):
@@
The optimal $\mathrm{AUGRC}$ is given by the second term in Equation 7, hence subtracting the optimal SC performance yields a pure ranking measure re-scaled by the probability of finding a positive/negative pair. This overcomes the lack of interpretability of $\mathrm{AURC}$ and $\mathrm{e\text{-}AURC}$ (R3). Monotonicity (R2) is ensured since the partial gradients w.r.t. both $\mathrm{AUROC_{f}}$ and $\mathrm{acc}$ are always negative. Further, it can accommodate arbitrary classification metrics via the error function $\ell$ (R4) as well as arbitrary CSF (R5).
@@ 4 Empirical Study
We conduct an empirical study based on the existing FD-Shifts benchmarking framework [Jäger et al., 2023], which compares various CSFs across a broad range of datasets. The focus of our study is not on the performance of individual methods (CSFs) but rather on the ranking of methods based on $\mathrm{AUGRC}$ evaluation as compared to $\mathrm{AURC}$. We utilize the same experimental setup as in Jäger et al. [2023] with the addition of CSF scores based on temperature-scaled classifier logits. A detailed overview of the datasets and methods used can be found in Appendix A.2. The code for reproducing our results and a PyTorch implementation of the $\mathrm{AUGRC}$ are available at: https://github.com/IML-DKFZ/fd-shifts.
@@ 4.1 Comparing Method Rankings of AUGRC and AURC
To study the relevance of $\mathrm{AUGRC}$, we investigate the changes induced by $\mathrm{AUGRC}$ on rankings of CSFs compared to $\mathrm{AURC}$ rankings.
CSF Rankings with AUGRC substantially deviate from AURC rankings. Figure 3 illustrates the ranking differences for all test datasets, showing changes in ranks across all CSFs. Notably, $\mathrm{AUGRC}$ induces changes in the top-3 methods (out of 13) on 5 out of 6 datasets. To ensure that these ranking changes are not due to variability in the test data, we evaluate the metrics on $500$ bootstrap samples from the test dataset and derive the compared method rankings based on the average rank across these samples ("rank-then-aggregate"). The size of each bootstrap sample corresponds to the size of the test dataset. Metric values for each bootstrap sample are averaged over 5 training initializations (2 for BREEDS and 10 for CAMELYON-17-Wilds). We analyze the robustness of the resulting method rankings for $\mathrm{AURC}$ and $\mathrm{AUGRC}$ separately, based on statistical significance. To that end, we perform pairwise CSF comparisons in form of a one-sided Wilcoxon signed-rank test [Wilcoxon, 1992] with a significance level of $\alpha=0.05$ on the bootstrap samples, as proposed in Wiesenfarth et al. [2021]. The resulting significance maps displayed in Figure 3 indicate stable rankings for both $\mathrm{AURC}$ and $\mathrm{AUGRC}$.This suggests that the observed ranking differences are induced by the conceptual difference of $\mathrm{AUGRC}$ compared to $\mathrm{AURC}$ and, as a consequence, that the shortcomings of current metrics and respective solutions discussed in Section 2, are not only conceptually sound, but also highly relevant in practice. For the results in Figure3, we select the DeepGamblers reward parameter and whether to use dropout for both metrics separately on the validation dataset (details in Table 1 and Table 2).
The shortcomings of AURC affect CSF comparison across datasets and distribution shifts. The method rankings for all datasets and distribution shifts, based on the original test datasets, are shown in Table 4, which also includes results for equal hyperparameters for $\mathrm{AURC}$ and $\mathrm{AUGRC}$. The substantial differences in method rankings across CSFs and datasets underline the relevance of $\mathrm{AUGRC}$ for SC evaluation. The complete $\mathrm{AUGRC}$ results on the FD-Shifts benchmark are shown in Table 3.
@@
Figure 3: Substantial differences in method rankings for AUGRC and AURC. On 5 out of 6 datasets, the top-3 CSFs (out of 13 compared CSFs) change when employing the proposed $\mathrm{AUGRC}$ instead of $\mathrm{AURC}$. This demonstrates the practical relevance of $\mathrm{AUGRC}$ for SC evaluation. CSFs are color-coded and sorted from top (best) to bottom (worst) by average rank based on $500$ bootstrap samples from the test dataset to ensure ranking stability. Ranking changes are reflected in changes in the color sequence and highlighted by red arrows. We assess the stability of the rankings for each metric individually using one-sided Wilcoxon signed-rank tests with a $5\%$ significance level based on the bootstrap samples. Adjacent to each ranking, we present the resulting significance maps for the pairwise CSF comparisons. These maps can be interpreted as follows: At each grid position $(x,y)$, filled entries indicate that metric values of CSF $y$ are ranked significantly better than those from CSF $x$ (across bootstrap samples), cross-marks indicate no significant superiority. An ideal ranking exhibits only filled entries above the diagonal.

### v2

@@ 1 Introduction
Several current approaches attempt to address this need for multi-threshold metrics in SC. The Area Under the Risk Coverage curve ($\mathrm{AURC}$) is the most prevalent of these Geifman et al. [2018]. However, we demonstrate in this work that $\mathrm{AURC}$ has significant limitations as it fails to adequately translate the risk from specific working points into a meaningful aggregated evaluation score. This shortcoming hampers the ability to benchmark and improve SC methodologies effectively.
@@
Proposal of AUGRC: We introduce the Area Under the Generalized Risk Coverage ($\mathrm{AUGRC}$) curve, a new metric designed to overcome the flaws of current multi-threshold metrics for SC. $\mathrm{AUGRC}$ meets all five requirements, providing a comprehensive and interpretable measure of SC system performance.
Empirical Validation: We empirically demonstrate the relevance and effectiveness of $\mathrm{AUGRC}$ through a comprehensive benchmark spanning 6 datasets and 13 confidence scoring functions.
@@
Figure 1: The AUGRC metric based on Generalized Risk overcomes common flaws in current evaluation of Selective classification (SC). a) Refined task definition for SC. Analogously to standard classification, we distinguish between holistic evaluation for method development and benchmarking using multi-threshold metrics versus evaluation of specific application scenarios at pre-determined working points. The current most prevalent multi-threshold metric in SC, $\mathrm{AURC}$, is based on Selective Risk, a concept for working point evaluation that is not suitable for aggregation over rejection thresholds (red arrow). To fill this gap, we formulate the new concept of Generalized Risk and a corresponding metric, $\mathrm{AUGRC}$ (green arrow). b) We formalize our perspective on SC evaluation by identifying five key requirements for multi-threshold metrics and analyze how previous metrics fail to fulfill them. Abbreviations, CSF: Confidence Scoring Function, AU(G)RC: Area Under the (Generalized) Risk Coverage curve.
@@ 2 Refined Task Formulation
Selective Classification (SC) systems consist of a classifier $m:\mathcal{X}\rightarrow\mathcal{Y}$, which outputs a prediction and a confidence scoring function (CSF) $g:\mathcal{X}\rightarrow\mathbb{R}$, which outputs a confidence score associated with the prediction. Assuming a supervised training setup, let $\{(x_{i},y_{i})\}_{i=1}^{N}$ be a dataset containing $N$ independent samples from the source distribution $P(X,Y)$ over $\mathcal{X}\times\mathcal{Y}$. Given a rejection threshold $\tau$, the model prediction is accepted only if the corresponding score is larger than $\tau$:
@@ 2.1 Evaluating SC systems in applied settings
Current Selective Classification evaluation commonly reports the selective risk of the system at a given coverage (“risk@coverage”), which implies a pre-determined cutoff on the risk or coverage (e.g. "maximum risk of 5%") leading to a working point of the system defined by a rejection threshold $\tau$. The selective risk is defined as:
@@
This formulation only considers accepted predictions with $c>\tau$. For a binary failure error, this risk is an empirical estimator for the conditional probability $P(Y_{\text{f}}=1|g(x)\geq\tau)$. Thus, this metric effectively communicates the risk of a "silent" failure for a specific prediction of classifier $m$ that has been accepted ($c>\tau$) at a pre-defined threshold $\tau$. This information may be useful in applied medical settings, e.g. to communicate the risk of misclassification for a patient p whose associated prediction has been accepted by a given SC system ($c_{p}>\tau$). Aside from the selective risk, other performance measurements for working point selection include for example Classification quality and Rejection quality [Condessa et al., 2017].
@@ 2.2 Evaluating SC systems for method development and benchmarking
Evaluating the general performance of a Selective Classification system requires a multi-threshold metric for a more comprehensive view compared to a working point analysis based on a fixed singular rejection threshold, analogous to binary classification tasks, where the Area Under the Receiver Operating Characteristic (AUROC) is used to assess the entire system performance across multiple classifier thresholds. However, the working point analysis can not be simply translated to multi-threshold aggregation, as it is based on Selective Risk, which only considers the risk w.r.t accepted predictions, assuming that a specific selection has already occurred ($P(Y_{\text{f}}=1\mid g(x)\geq\tau)$). This evaluation ignores rejected cases and thus contradicts the holistic assessment, where we are interested in the general risk for any prediction causing a silent failure independent of the future selection decision of individual predictions. This holistic assessment is reflected by the joint probability $P(Y_{\text{f}}=1,g(x)\geq\tau)$, which reflects the risk of silent failure for any prediction processed by the system including cases that are potentially rejected in the future.
@@ 2.3 Requirements for Selective Classification multi-threshold metrics
Based on the refined task definition given above, we can formulate concrete five requirements R1-R5 for multi-threshold metrics in Selective Classification:
R1: Task Alignment. The general goal in SC is to prevent silent failures either by preventing failures in the first place (via classifier performance), or by detecting the remaining failures (via CSF ranking quality). As argued in Jäger et al. [2023], it is crucial to jointly evaluate both aspects of an SC system, since the choice of CSF generally affects the underlying classifier. R2: Monotonicity. The metric should be monotonic w.r.t both evaluated factors stated in R1, i.e. improving on one of the two factors while keeping the other one fixed results in an improved metric value. Note that R2 does not make assumptions about how the two factors are combined, but it represents a minimum requirement for meaningful comparison and optimization of the metric. R3: Ranking Interpretability. Interpretability is a crucial component of a metric [Maier-Hein et al., 2022]. AUROC is the de facto standard ranking metric for binary classification tasks, providing an intuitive assessment of ranking quality which is proportional to the number of permutations needed for establishing an optimal ranking. This can also be interpreted as the "probability of a positive sample having a higher score than a negative one". Ideally, the SC metric should follow this intuitive assessment of rankings. R4: CSF Flexibility. The metric should be applicable to arbitrary choices of CSFs. This includes external CSFs, i.e. scores that are not based directly on the classifier output. R5: Error Flexibility. Current SC literature largely focuses on 0/1 error (i.e. $1-\text{accuracy}$) in their risk computation. However, in many real-world scenarios, accuracy is not an adequate classification metric, such as in the presence of class imbalance and for pixel-level classification. Thus, the SC metric should be flexible w.r.t the choice of error function.
@@ 2.4 Current multi-threshold metrics in SC do not fulfill requirements R1-R5
This integral effectively aggregates the Selective Risk (Equation 2) over all fractions of accepted predictions (i.e. coverage). However, as discussed in Section 2.2, the Selective Risk is not suitable for aggregation over rejection thresholds to holistically assess a SC system. This inadequacy effectively leads to an excessive over-weighting of high-confidence failures in $\mathrm{AURC}$ and in return to breaching the requirements of monotonicity (R2) and ranking interpretability (R3). We provide empirical evidence for these shortcomings on toy data (Section 3) and real-world data (Section 4.2). Further, current SC literature usually employs the 0/1 error function, even though the related Accuracy metric is well known to be an unsuitable classification performance measure for many applications. For example, Balanced Accuracy is used to address class imbalance and metrics such as the Dice Score are employed for segmentation. Moving beyond the 0/1 error yields a higher variance in distinct error values and may thus amplify the shortcoming of over-weighting high-confidence failures. As more sophisticated error functions being are an important direction of future SC research, it is crucial to ensure the compatibility of metrics in SC evaluation.
e-AURC: Geifman et al. [2018] further introduce the excess-$\mathrm{AURC}$ as the difference to the $\mathrm{AURC}$ of an optimal confidence scoring function (denoted as $\mathrm{AURC}^{*}$)
@@ 3 Area under the Generalized Risk Coverage Curve
This metric evaluates SC models in terms of the rate of silent failures averaged over working points and thus provides a practical measurement that is directly interpretable. It is bounded to the $[0,\nicefrac{{1}}{{2}}]$ interval, whereby lower $\mathrm{AUGRC}$ corresponds to better SC performance. The $\mathrm{AUGRC}$ is not subject to the shortcomings of the $\mathrm{AURC}$, and we can derive a direct relationship to $\mathrm{AUROC_{f}}$ and $\mathrm{acc}$ (the derivation and visualizations are shown in Equation 8 and Figure 5):
@@
To illustrate, when drawing two random samples, it corresponds to the probability that either both are failure cases or that one is a failure and has a higher confidence score than the non-failure. This is reflected by the second and first term of Equation 7, respectively. As an example, a CSF that outputs random scores yields $\mathrm{AUROC_{f}}=\frac{1}{2}$, and hence $\mathrm{AUGRC}=\frac{1}{2}(1-\mathrm{acc})$. The $\mathrm{AUGRC}$ for an optimal CSF ($\mathrm{AUROC_{f}}=1$) is given by the second term in Equation 7, hence subtracting the optimal SC performance yields a pure ranking measure re-scaled by the probability of finding a positive/negative pair. This overcomes the lack of interpretability of $\mathrm{AURC}$ and $\mathrm{e\text{-}AURC}$ (R3). Monotonicity (R2) is ensured since the partial gradients w.r.t. both $\mathrm{AUROC_{f}}$ and $\mathrm{acc}$ are always negative. Further, it can accommodate arbitrary classification metrics via the error function $\ell$ (R4) as well as arbitrary CSF (R5).
@@ 4 Empirical Study
We conduct an empirical study based on the existing FD-Shifts benchmarking framework [Jäger et al., 2023], which compares various CSFs across a broad range of datasets. The focus of our study is not on the performance of individual methods (CSFs) but rather on the ranking of methods based on $\mathrm{AUGRC}$ evaluation as compared to $\mathrm{AURC}$. We utilize the same experimental setup as in Jäger et al. [2023] with the addition of CSF scores based on temperature-scaled classifier logits. Extending the benchmark by recent SC methods, e.g. Feng et al. [2022], is an interesting direction of future work. A detailed overview of the datasets and methods used can be found in Appendix A.2. The code for reproducing our results and a PyTorch implementation of the $\mathrm{AUGRC}$ are available at: https://github.com/IML-DKFZ/fd-shifts.
@@ 4.1 Comparing Method Rankings of AUGRC and AURC
To study the relevance of $\mathrm{AUGRC}$, we investigate the changes induced by $\mathrm{AUGRC}$ on rankings of CSFs compared to $\mathrm{AURC}$ method rankings.
CSF Rankings with AUGRC substantially deviate from AURC rankings. Figure 3 illustrates the method ranking differences for all i.i.d. test datasets, showing changes in ranks across all CSFs. Notably, $\mathrm{AUGRC}$ induces changes in the top-3 methods (out of 13) on 5 out of 6 datasets. To ensure that these ranking changes are not due to variability in the test data, we evaluate the metrics on $500$ bootstrap samples from the test dataset and derive the compared method rankings based on the average rank across these samples ("rank-then-aggregate"). The size of each bootstrap sample corresponds to the size of the test dataset. Metric values for each bootstrap sample are averaged over 5 training initializations (2 for BREEDS and 10 for CAMELYON-17-Wilds). We analyze the robustness of the resulting method rankings for $\mathrm{AURC}$ and $\mathrm{AUGRC}$ separately, based on statistical significance. To that end, we perform pairwise CSF comparisons in form of a one-sided Wilcoxon signed-rank test [Wilcoxon, 1992] on the bootstrap samples. To control the family-wise error rate at a 5% significance level, we apply the Holm correction for multiple testing per metric and dataset, following [Wiesenfarth et al., 2021] (see Figure 7 for the results without correction for multiple testing). The resulting significance maps displayed in Figure 3 indicate stable method rankings for both $\mathrm{AURC}$ and $\mathrm{AUGRC}$. This suggests that the observed ranking differences are induced by the conceptual difference of $\mathrm{AUGRC}$ compared to $\mathrm{AURC}$ and, as a consequence, that the shortcomings of current metrics and respective solutions discussed in Section 2, are not only conceptually sound, but also highly relevant in practice. For the results in Figure 3, we select the DeepGamblers reward parameter and whether to use dropout for both metrics separately on the validation dataset (details in Table 1 and Table 2).
The shortcomings of AURC affect CSF comparison across datasets and distribution shifts. The method rankings for all datasets and distribution shifts, based on the original test datasets, are shown in Table 4, which also includes results for equal hyperparameters for $\mathrm{AURC}$ and $\mathrm{AUGRC}$. Results for the bootstrap-based method ranking analysis for distribution shifts are displayed in Figure 6. The substantial differences in method rankings across all datasets and distribution shifts underline the relevance of $\mathrm{AUGRC}$ for SC evaluation. The complete $\mathrm{AUGRC}$ results on the FD-Shifts benchmark are shown in Table 3.
@@
Figure 3: Substantial differences in method rankings for AUGRC and AURC. On 5 out of 6 datasets, the top-3 CSFs (out of 13 compared methods) change when employing the proposed $\mathrm{AUGRC}$ instead of $\mathrm{AURC}$. This demonstrates the practical relevance of the $\mathrm{AUGRC}$ metric for Selective Classification evaluation. CSFs are color-coded and sorted from top (best) to bottom (worst) by average rank based on $500$ bootstrap samples from the test dataset to ensure ranking stability. Ranking changes are reflected in changes in the color sequence and highlighted by red arrows. We assess the stability of the method rankings for each metric individually using one-sided Wilcoxon signed-rank tests based on the bootstrap samples at $5\%$ significance level with adjustment for multiple testing according to Holm. Adjacent to each ranking, we present the resulting significance maps for the pairwise CSF comparisons. These maps can be interpreted as follows: At each grid position $(x,y)$, filled entries indicate that metric values of CSF $y$ are ranked significantly better than those from CSF $x$ (across bootstrap samples), cross-marks indicate no significant superiority. An ideal ranking exhibits only filled entries above the diagonal.